M and M, Inc. produces a product that has a variable cost of $4.20 per unit. The companys fixed costs are $50,400. The product is sold for $7 per unit and the company desires to earn a target profit of $11,200. What is the amount of sales that will be necessary to earn the desired profit? (Do not round intermediate calculations)
Answer #1
Unit Contribution margin = 7-4.20 = $2.80
Contribution margin ratio = 2.80/7= 40%
Amount of sales required = (Fixed costs+Target Profit)/Contribution margin ratio = (50400+11200)/40%= $154000